OpenStudy (anonymous):

Which is the best system of equation to use for the following equations: 6x – 4y = 54 –9x + 2y = –69

OpenStudy (ash2326):

elimination, multiply second by 2, and add it to first to eliminate y then find x---> use it to find y

hero (hero):

You're still doing these?

OpenStudy (anonymous):

Can you guys explain with exact steps?

hero (hero):

You don't know how to multiply the second equation by 2?

OpenStudy (ash2326):

6x – 4y = 54 –9x + 2y = –69 multiply equation 2 by 2 -18x+4y = -138 add this to first 6x-4y-18x+4y=54-138 -12x=-84 x=7 put x in first 6*7-4y=54 42-4y=54 -4y=12 y=-3 so x=7 and y=-3